a b and c share 7560 naira so that a has 2 and a half times as much as c, and b has 3 and a half times as much as c. find their shares
Answer by mananth(16946)

a b and c share 7560 naira so that a has 2 and a half times as much as c, and b has 3 and a half times as much as c. find their shares
a= 2.5c
b= 3.5c

total naira =7560
a+b+c=7560
2.5c+3.5c+c =7560
7c = 7560
c= 7560 /7
c= 1080
c gets 1080
b= 2.5 c
a= 2.5*1080 = 2700
b=3.5*1080= 3780
CHECK
a+b+c = 1080 +2700 +3780= 7560